# Few-Shot Learner Generalizes Across AI-Generated Image Detection

### v1

@@ 3.3 Zero-shot Classification
We propose a zero-shot detection way to verify the applicability of FSD in real-world settings. In the zero-shot task, there is no additional samples from the test classes. Instead, we are provided with a meta-data vector for each class in the training set to categorize the unseen images. We place considerable emphasis on zero-shot classification, since in practical scenarios, it is unfeasible to obtain test samples from every possible generative model. The zero-shot results can represent the performance of our model in general scenarios.
In this work, we simply define the meta-data vector as the prototypical representation derived from an extensive collection of samples. That is, we randomly select plenty of images from each class in the training set and then compute their prototypical representations following Equation 1. Once the nearest representation of a test image is labeled as a synthetic class, it will be regarded as a fake image.
4 Experiment
Table 1: Comparison of our FSD and previous studies. We train several classifiers on different subsets and report the accuracy for each unseen test subset. The best results are highlighted in boldface, while the second-optimal results are marked with underline.
@@ 4 Experiment
CNNSpot (Wang et al., 2020) 58.2 57.1 57.0 70.3 56.7 56.6 59.3
Spec (Zhang et al., 2019b) 56.7 65.4 57.9 72.4 61.7 64.3 63.1
F3Net (Qian et al., 2020) 55.1 57.8 66.5 73.1 62.1 56.5 61.9
GramNet (Liu et al., 2020) 58.1 65.3 58.7 72.8 57.8 61.2 62.3
DIRE (Wang et al., 2023) 65.0 69.1 61.9 73.7 63.4 56.7 65.0
LARE2 (Luo et al., 2024) 66.4 81.3 66.7 87.3 84.4 74.0 76.7
@@ 4.1 Datasets and Evaluation Metrics
Datasets. We evaluate our proposed method on the widely used GenImage dataset (Zhu et al., 2023), which contains $1,331,167$ real images from ImageNet (Deng et al., 2009) and $1,350,000$ generated images from $7$ diffusion models and one GAN. The diffusion generators include Midjourney (mid, 2022), Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), Wukong (wuk, 2022), ADM (Dhariwal & Nichol, 2021), GLIDE (Nichol et al., 2022) and VQDM (Gu et al., 2022). And the only GAN is BigGAN (Brock et al., 2019). The real images are first divided into $8$ subsets. And each subset is then further partitioned into training part and test part. Each generator is paired with a specific subset and uses the category labels of the real images within the subset to produce synthetic images. Therefore, GenImage dataset consists of $8$ fake-vs-real subsets corresponding to the $8$ generative models.
In this study, the synthetic images in different subsets are categorized into distinct classes, whereas the real images across all subsets are classified into a single class. We also notice that $3$ diffusion models SD v1.4, SD v1.5 and Wukong share an identical model structure and it is difficult to distinguish among them. Thus, we combine their corresponding subsets into a unified subset, named Stable Diffusion (SD). Finally, the dataset used in our experiments comprises $7$ classes of images from different sources.
@@ 4.2 Implementation Details
To be more comparable with previous works, we adopt the ResNet-$50$ (He et al., 2016) pretrained on ImageNet (Deng et al., 2009) as the backbone of our model, which outputs a prototype vector of 1024 dimensions. Following Wang et al. (2023), the input images are first resized to $256\times 256$ and then randomly cropped to $224\times 224$ with a random horizontal flip during training. In contrast, only a center crop to $224\times 224$ is performed after resizing the images during testing. The distance in the metric space is measured with squared Euclidean distance.
To fit the classification task, we apply different few-shot manners to training and testing. In detail, we select one synthetic subset as the test data and utilize the rest $6$ subsets as the training data for one classifier. At every training step, $3$ classes are randomly selected from the training set, with $5$ random samples chosen from each class for the support batch and another $5$ samples for the query batch. We employ Adam as the optimizer to minimize the cross entropy loss with a base learning rate of $10^{-4}$. We also adopt a StepLR scheduler with $\gamma=0.5$ and $step=80000$. The training of each classifier is carried out for $200,000$ steps with a batch size of $16$. During testing, we conduct a $2$-way classification task by comparing the selected test class against the real class. For the zero-shot detection, we randomly select $1024$ samples from each class in the training set to compute the meta-data vectors. Our method is implemented by PyTorch library and all the experiments are conducted on a single A100 with 40GB memory.
Table 2: Result of $10$-shot cross-generator image classification on different subsets. Each classifier is trained on the dataset excluding the subset from the first column. All classifiers are tested on the $6$ synthetic classes, reporting Accuracy / Average Precision (%).
@@ 4.3 Comparison to Existing Methods
We compare FSD with several state-of-the-art synthetic image detection methods, including CNNSpot (Wang et al., 2020), Spec (Zhang et al., 2019b), F3Net (Qian et al., 2020), GramNet (Liu et al., 2020), DIRE (Wang et al., 2023) and LARE2 (Luo et al., 2024). The first four methods are model-agnostic, focusing on identifying synthetic artifacts within fake images. And the last two methods leverage the reconstruction error of Stable Diffusion to differentiate between fake and real images. To conduct a comprehensive comparison, we train $6$ classifiers on different training sets, each of which excludes a specific test subset. However, previous studies typically provide several classifiers, and each classifier is trained on one distinct subset and evaluated on the others. To minimize the difference between our method and previous researches, we report the average performance of those classifiers over each test subset in previous studies, which can better represent their generalization ability on each class. The results are summarized in Table 1.
As illustrated in Table 1, our FSD attains state-of-the-art performance across $5$ out of the $6$ test classes, surpassing previous state-of-the-art by $+7.4\%$ accuracy in average. The suboptimal performance of our model on the VQDM class can be attributed to the involvement of image quantization in VQDM, which significantly differs from the other generative models in our experiments. Another reason is that LARE2 is specifically designed to detect synthetic images generated by diffusion models, which enables it to achieve unexpectedly high performance on the VQDM class. Although our few-shot method may not be directly comparable to other approaches due to the accessing of the few samples in the test class, we still regard this discrepancy as a significant strength of our model. The definite improvement in performance demonstrates FSD’s capability to effectively leverage a limited number of unseen samples.
Based on Table 1, we find that FSD achieves notable performance in zero-shot scenarios, in which the prototypical representations are calculated from samples in the training classes. The result shows that our approach is also suitable for detecting synthetic images even when samples from the same sources are not provided, thus broadening the application scope of FSD. We owe this advantage to FSD’s capability to capture not only the intra-class characteristics but also the common features across all classes. Another observation is that the discrepancy between zero-shot results and few-shot results remains notable, which indicates the necessity of gathering training data from a wider range of generative models.
@@ 4.4 Cross-Generator Classification
As shown in Table 2, each row presents the classification results for one classifier across all the $6$ synthetic classes. The results show that it is easy for a classifier to detect fake images from those classes encountered during training, which is consistent with the previous studies (Zhu et al., 2023; Luo et al., 2024). When it comes to detecting the unseen classes, as shown in the diagonal of Table 2, FSD suffers from a slight decline in performance, demonstrating the generalization capability of our model. Another finding is that FSD performs nearly perfectly on the VQDM class when this class has been included in the training set. However, the performance significantly declines when this class has been excluded. This fact shows that it is still challenging to generalize across unseen generative models, indicating that our approach remains inadequate in capturing the discriminative features between classes.
We also observe the varying performance of FSD across different classes. For instance, images generated by GLIDE, ADM, SD and VQDM are relatively easy to detect by our model when these classes are included in the training set, while those from Midjourney and BigGAN are a little challenging to identify. Particularly, images from the GLIDE class can be distinguished with exceptional accuracy, even when not included in the training set. These facts indicate the necessity of treating images from different sources as distinct categories, which has not been fully explored in previous researches. Moreover, it is essential to collect images from representative generative models for training, as generalization on similar generators is easier than on those that are vastly different.
@@ 4.5 Visualization
To further analyze the effectiveness of FSD, we utilize t-SNE (van der Maaten & Hinton, 2008) visualization to illustrate the feature space of classifiers on different training sets, as shown in Figure 3. Each classifier is trained on those subsets excluding the test one in each sub-figure. We randomly selected $1024$ samples from the test part of each class for visualization. And the features are extracted from the final layer of our model.
From Figure 3, we can see that features of images from different generative models exhibit distinct distributions, each of which is marked by a unique color. Most classes within the training set can be well-separated due to the large margins among their distributions. However, the boundary of the Midjourney class represented in red is not as clear as that of the others. This can be attributed to its API service, which may utilize multiple generative models for image generation based on our inference. Although the boundary of the test class in each sub-figure is a little ambiguous, the distribution of this class still remains identifiable, which demonstrates that our method successfully captures the intra-category features of unseen data.
The results also support the validity of our approach that categorizes samples from different generators as different classes. However, such categorization also introduces the issue of eliminating the significant differences between seen data and unseen data. As illustrated in Figure 3(e), The distribution of the VQDM class in orange exhibits a large overlap with that of the real class in gray, due to the great difference in image tokenization between VQDM and other generative models. This also indicates that our method still faces the challenge in detecting images from models that are significantly different from those generative models used in training, which will be a main focus of our future work.
@@ 4.6 Ablation Study
Impact of the number of shots. An important aspect is to determine the minimum number of samples required by our approach to achieve satisfactory performance. We choose different numbers of support samples in each class, ranging from $1$ to $200$. To present a comprehensive overview of this aspect, we conducted the experiment across all the $6$ classifers described in Section 4.4. We evaluate each classifier on the corresponding test class with varying numbers of support samples. To maintain a consistent total number of test samples across different settings, We set a fixed ratio of $1:3$ between the support set and the query set. The results are shown in Figure 4, in which each fold line represents the performance of a classifier across varying sizes of support set.
From Figure 4, we can see that using more support samples results in better performance. However, it also increases the required computational resources. We find that detecting with $10$ shots achieves an optimal balance between performance and resource consumption for most of the test classes. For example, our model brings the accuracy on the ADM class from $62.6\%$ to $79.2\%$ when increasing the number of shots from $1$ to $10$, with an improvement of $+16.6\%$. However, it achieves $81.7\%$ accuracy when using $200$ shots, gaining only an increase of $+2.5\%$ over the $10$-shots scenario. As it is easy to collect $10$ samples from online generative models and the computation consumption is very little, We ultimately decide to use the $10$-shots results of FSD for comparison.
@@
Figure 4: Influences of the number of shots. Each classifier is trained on those subsets excluding the test one. And $10$-shots is the most cost-effective setting across all the categories.
@@
Figure 5: Visualization of network features for our FSD (a) and a binary classifier (b) on ADM class. In the feature space of FSD, images from the unseen class in purple are tightly clustered together. However, they are much more dispersed in the feature space of the binary classifier.
Impact of the multi-class classification. Different from previous studies which treat the fake images as a single class and train an binary classifier to detect them, we category the images from different generators as distinct classes. We argue that distinguishing images from different generators can also benefit the detection performance. Thus, we train another binary classifier and compare it with our method on the ADM class, which is difficult to detect according to our experiments. Both classifiers utilize the ResNet-$50$ as backbone and are trained on the dataset excluding the ADM class. We also apply the t-SNE visualization and random select $1024$ samples for visualizing. Features of the binary classifier are extracted from the output layer of the ResNet-$50$, with a dimension of $2048$. The visualization results are shown in Figure 5.
Figure 5(a) shows that our FSD successfully distinguishes those images across different classes which has been used in training. Images from the unseen ADM class dyed in purple also cluster together in the visualization picture. In contrast, the traditional binary classifier fails to distinguish images from Midjourney and Stable Diffsion, as shown in Figure 5(b). And the unseen samples in purple are quite dispersed, failing to form a compact cluster. Thus, distinguishing them from the real ones can be extremely difficult. This fact indicates that our method is able to extract intra-category common features from unseen classes, thereby enabling an accurate differentiation among various classes. Moreover, we find that it is difficult for a classifier to distinguish samples from Stable Diffusion V1.4 and Stable Diffusion V1.5. One plausible explanation could be the shared usage of an identical backbone and VAE, which makes their products significantly similar. This implies that we should carefully select the representative generative models to build the training set, which We leave for future work.
@@ 5 Conclusion and Limitations
The key point of detecting fake images in real-world scenarios is to make full use of the limited images from unseen generative models. In this paper, we propose Few-Shot Detector (FSD), a novel AI-generated image detector based on few-shot learning. Experiments show that FSD achieves state-of-the-art performance in synthetic image detection with only $10$ samples from the test generator, demonstrating the strong generalization ability of our model. We visualize the feature space to highlight the advantage of distinguishing images from different sources, which has not been widely adopted among previous studies.
However, our method still suffers from a few limitations. Since it is not feasible to obtain few-shot samples from all generative models, our collection of synthetic images is confined to part of them, which may lead to a decrease in detection performance. Another weakness is that our method does not perform satisfactorily when detecting images from a vastly different generative model. We aim to address these issues in our future work.

### v2

@@ 3.3 Zero-shot Classification
We propose a zero-shot detection approach to verify the applicability of FSD in real-world settings. In the zero-shot task, there are no additional samples from the test classes. Instead, we are provided with a metadata vector for each class in the training set to categorize the test images. We put considerable emphasis on this task, since acquiring samples from all potential generative models can be impractical. The zero-shot results can represent the performance of our model in general scenarios.
Table 1: Comparison of our FSD with previous studies. We train several classifiers on different subsets and report the accuracy for each unseen test subset. The best results are highlighted in boldface, while the second-optimal results are marked with underline.
@@
Spec (Zhang et al., 2019b) 50.0 64.7 52.8 56.1 56.5 63.0 57.2
CNNSpot (Wang et al., 2020) 52.8 73.3 55.0 55.9 54.4 66.2 59.6
F3Net (Qian et al., 2020) 50.1 52.5 66.4 57.0 59.4 50.4 56.0
GramNet (Liu et al., 2020) 51.3 62.6 53.8 56.8 52.2 57.2 55.7
DIRE (Wang et al., 2023) 57.9 68.2 57.3 58.2 59.6 50.8 58.7
LARE2 (Luo et al., 2024) 62.7 80.2 63.5 79.6 76.9 72.0 72.5
@@
In this work, we simply define the metadata vector as the prototypical representation derived from an extensive collection of samples. That is, we randomly select a large number of images from each class in the training set and then compute the prototypical representations following Equation 1. If the nearest representation of a test image is labeled as a synthetic class, the image will be considered fake.
4 Experiment
@@ 4.1 Datasets and Evaluation Metrics
Datasets. We evaluate our proposed method on the widely used GenImage dataset (Zhu et al., 2023), which contains $1,331,167$ real images from ImageNet (Deng et al., 2009) and $1,350,000$ generated images from $7$ diffusion models and one GAN. The diffusion generators include Midjourney (mid, 2022), Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), Wukong (wuk, 2022), ADM (Dhariwal & Nichol, 2021), GLIDE (Nichol et al., 2022) and VQDM (Gu et al., 2022). The only GAN utilized is BigGAN (Brock et al., 2019). The real images are first divided into $8$ subsets. Each subset is subsequently partitioned into a training part and a test part. Each generator is associated with one specific subset and uses the category labels of the real images within the subset to produce synthetic images. Consequently, the GenImage dataset consists of $8$ fake-vs-real subsets for analysis.
In this study, the synthetic images in different subsets are categorized into distinct classes, while real images from all subsets are aggregated into a single class. We also observe that $3$ diffusion models, SD v1.4, SD v1.5 and Wukong, share an identical model structure, making it challenging to differentiate among them. Thus, we merge their corresponding subsets into a unified subset, named Stable Diffusion (SD). Ultimately, the dataset used in our experiments comprises $7$ classes of images from different sources.
@@ 4.2 Implementation Details
To be more comparable with previous works, we adopt the ResNet-$50$ (He et al., 2016) pretrained on ImageNet (Deng et al., 2009) as the backbone of our model, which outputs a prototype vector of 1024 dimensions. Following Wang et al. (2023), the input images are first resized to $256\times 256$ and then randomly cropped to $224\times 224$ with random horizontal flipping during training. In contrast, only a center crop to $224\times 224$ is performed after resizing the images during testing. The distance in the metric space is measured with Squared Euclidean Distance.
To fit the classification task, we employ different few-shot strategies for training and testing. Specifically, we select one synthetic subset as the test set and utilize the remaining $6$ subsets to train a classifier. At every training step, $3$ classes are randomly selected from the training set, with $5$ samples chosen from each class for the support set and another $5$ samples per class for the query set. We employ Adam as the optimizer to minimize the cross-entropy loss with a base learning rate of $10^{-4}$. We also adopt a StepLR scheduler with $\gamma=0.5$ and $step\_size=80000$. Each classifier is trained for $200,000$ steps, with a batch size of $16$. During testing, we conduct a binary classification task by comparing the selected test class against the real class. For zero-shot detection, we randomly select $1024$ samples from each training subset to compute the metadata vectors. Our method is implemented with the PyTorch library and all the experiments are conducted on a single A100 with 40GB memory.
Table 2: Results of $10$-shot cross-generator image classification on different subsets. Each classifier is trained on the dataset excluding the subset from the first column. All classifiers are tested on the $6$ synthetic classes, reporting accuracy/average precision (%).
@@ 4.3 Comparison to Existing Methods
We compare FSD with several state-of-the-art synthetic image detection methods and summarize the results in Table 1. The compared group includes CNNSpot (Wang et al., 2020), Spec (Zhang et al., 2019b), F3Net (Qian et al., 2020), GramNet (Liu et al., 2020), DIRE (Wang et al., 2023) and LARE2 (Luo et al., 2024). The first four methods are model-agnostic, focusing on identifying synthetic artifacts within fake images. The last two methods leverage the reconstruction error of Stable Diffusion to differentiate between fake and real images. To conduct a comprehensive comparison, we train $6$ classifiers on different training sets, each of which excludes a specific test subset. However, previous studies typically provide several classifiers, and each classifier is trained on one distinct subset and evaluated on the others. To minimize the difference between our method and previous works, we report the average performance of $5$ classifiers trained on non-test categories and evaluated on each test subset, which can better represent their generalization ability on each class.
As illustrated in Table 1, our FSD attains state-of-the-art performance across $5$ out of the $6$ test classes, surpassing previous state-of-the-art by $+11.6\%$ accuracy on average. The suboptimal performance of our model on the VQDM class can be attributed to the involvement of image quantization in VQDM, which significantly differs from the other generative models in our experiments. Another reason is that LARE2 is specifically designed to detect synthetic images generated by diffusion models, which enables it to achieve unexpectedly high performance on the VQDM class. Although our few-shot method may not be directly comparable to other approaches due to accessing only a few samples in the test class, we still regard this discrepancy as a significant strength of our model. The definite improvement in performance demonstrates FSD’s capability to effectively leverage a limited number of unseen samples.
Based on Table 1, we find that FSD achieves notable performance in zero-shot scenarios, in which the prototypical representations are calculated from samples in the training classes. These results show that our approach is also suitable for detecting synthetic images even when samples from the same sources are not provided, thus broadening the application scope of FSD. We attribute this advantage to FSD’s ability to capture not only the intra-class characteristics but also the common features across all classes. Another observation is that the discrepancy between zero-shot results and few-shot results remains notable, which indicates the necessity of gathering training data from a wider range of generative models.
@@ 4.4 Cross-generator Classification
As shown in Table 2, each row presents the classification results for one classifier across all the $6$ synthetic classes. The results show that it is easy for a classifier to detect fake images from those classes encountered during training, which is consistent with the previous studies (Zhu et al., 2023; Luo et al., 2024). When it comes to detecting the unseen classes, as shown in the diagonal of Table 2, FSD suffers from a slight decline in performance, demonstrating the generalization capability of our model. Another finding is that FSD performs nearly perfectly on the VQDM class when this class is included in the training set. However, the performance significantly declines when this class has been excluded. This fact shows that it is still challenging to generalize across vastly different unseen generative models, indicating that our approach remains inadequate in capturing the discriminative features between classes.
@@
We also observe the varying performance of FSD across different classes. For instance, images generated by GLIDE, ADM, SD and VQDM are relatively easy to detect by our model when these classes are included in the training set, while those from Midjourney and BigGAN are a little challenging to identify. Particularly, images from the GLIDE class can be distinguished with exceptional accuracy, even when not included in the training set. These facts indicate the necessity of treating images from different sources as distinct categories, which has not been fully explored in previous research. Moreover, it is essential to collect images from representative generative models for training, as generalization on similar generators is easier than on those that are vastly different.
@@ 4.5 Visualization
To further analyze the effectiveness of FSD, we utilize t-SNE (van der Maaten & Hinton, 2008) visualization to illustrate the feature space of classifiers on different training sets, as shown in Figure 4. Each classifier is trained on those subsets excluding the test one in each sub-figure. We randomly selected $1024$ samples from the test part of each class for visualization. The features are extracted from the final layer of our model.
From Figure 4, we can see that features of images from different generative models exhibit distinct distributions, each of which is marked by a unique color. Most classes within the training set can be well-separated due to the large margins among their distributions. However, the boundary of the Midjourney class represented in red is not as clear as that of the others. This can be attributed to its API service, which may utilize multiple generative models for image generation as inferred from our analysis. Although the boundary of the test class in each sub-figure is a little ambiguous, the distribution of this class still remains identifiable, which demonstrates that our method successfully captures the intra-category features of unseen data.
The results also support the validity of our approach that categorizes samples from different generators as different classes. However, such categorization also introduces the issue of eliminating the significant differences between seen data and unseen data. As illustrated in Figure 4(e), the distribution of the VQDM class in orange exhibits a large overlap with that of the real class in gray, due to VQDM’s significant differences in image tokenization compared to other models. This also indicates that our method still faces the challenge of detecting images from models that are significantly different from those generative models used in training, which will be the main focus of our future work.
@@ 4.6 Ablation Study
Impact of the number of shots. An important aspect is to determine the minimum number of samples required by our approach to achieve satisfactory performance. We choose different numbers of support samples in each class, ranging from $1$ to $200$. To present a comprehensive overview of this aspect, we conduct the experiment across all the $6$ classifiers described in Section 4.4. We evaluate each classifier on the corresponding test class with varying numbers of support samples. To maintain a consistent total number of test samples across different settings, we set a fixed ratio of $1:3$ between the support set and the query set. The results are shown in Figure 5. Each line represents the performance of one classifier across varying sizes of support set.
From Figure 5, we can see that using more support samples results in better performance. However, it also increases the required computational resources. We find that detecting with $10$ shots achieves an optimal balance between performance and resource consumption for most of the test classes. For example, our model brings the accuracy on the ADM class from $62.6\%$ to $79.2\%$ when increasing the number of shots from $1$ to $10$, with an improvement of $+16.6\%$. However, it achieves $81.7\%$ accuracy when using $200$ shots, gaining only an increase of $+2.5\%$ over the $10$-shots scenario. As it is easy to collect $10$ samples from online generative models and the computational cost is very low, we ultimately decide to use the $10$-shots results of FSD for comparison.
@@
Figure 5: Influence of the number of shots. Each classifier is trained on those subsets excluding the test one. The $10$-shot setting is the most cost-effective setting across all categories.
Impact of the multi-class classification. Different from previous studies that treat the fake images as a single class and train a binary classifier to detect them, we categorize the images from different generators as distinct classes. We argue that distinguishing images from different generators can also benefit the detection performance. Thus, we train another binary classifier and compare it with our method on the ADM class, which is difficult to distinguish according to our experiments. Both classifiers utilize the ResNet-$50$ as the backbone and both of them are trained on the dataset excluding the ADM class. We also apply the t-SNE visualization and randomly select $1024$ samples for visualizing. Features of the binary classifier are extracted from the output layer of the ResNet-$50$, with dimensions of $2048$. The visualization results are shown in Figure 6.
Figure 6(a) shows that our FSD successfully distinguishes those images across different classes which have been used in training. Images from the unseen ADM class colored in purple also cluster together in the visualization picture. In contrast, the traditional binary classifier fails to distinguish images from Midjourney and Stable Diffusion, as shown in Figure 6(b). The unseen samples in purple are quite dispersed, failing to form a compact cluster. Thus, distinguishing them from the real ones can be extremely difficult. This fact indicates that our method possesses a certain ability to extract intra-category commonality from unseen classes, thereby enabling an accurate differentiation among various classes. Moreover, we find that it is difficult for a classifier to distinguish samples from Stable Diffusion V1.4 and Stable Diffusion V1.5. One plausible explanation could be the shared usage of an identical backbone and VAE, which makes their products highly similar. This implies that we should carefully select the representative generative models to build the training set, which we leave for future work.
@@
Figure 6: Visualization of network features for our FSD (a) and a binary classifier (b) on the ADM class. In the feature space of FSD, images from the unseen class in purple are tightly clustered together. However, they are significantly more dispersed in the feature space of the binary classifier.
@@ 5 Conclusion and Limitations
The key point of detecting fake images in real-world scenarios is to make full use of the limited number of images from unseen generative models. In this paper, we propose the Few-Shot Detector (FSD), a novel AI-generated image detector based on few-shot learning. Experiments show that FSD achieves state-of-the-art performance in synthetic image detection with only $10$ samples from the test generators, demonstrating the strong generalization ability of our model. We visualize the feature space to highlight the advantage of categorizing synthetic images by their sources, an approach that has not been widely adopted in previous studies.
However, our method still suffers from a few limitations. Since it is not feasible to obtain few-shot samples from all generative models, our collection of synthetic images is confined to part of these models, which may lead to a decrease in detection performance. Another weakness is that our method does not perform satisfactorily when detecting images from a vastly different generative model. We aim to address these issues in our future work.